Kitty puts ¼ of her the money she earned from selling embutido in savings and uses ½ of the remaining money to pay back her father (who loaned her the capital). If she has Php 1500 left, how much did she have in the first place?
Answer by Theo(13342)

let x be the amount of money she had in the first place.

she put 1/4 * x into savings.

she had 3/4 * x left.

she gave 1/2 * 3/4 * x = 3/8 * x to her father.

she kept 1/2 * 3/4 * x = 3/8 * x for herself.

the half she kept for herself was worth 1500 php.

that means 3/8 * x = 1500

solve for x to get x = 8/3 * 1500 = 4000 php.

she had 4000 php to start.

she saved 1/4 * 4000 = 1000 php.

she had 3/4 * 4000 = 3000 php left.

she gave half to her father and saved half = 1500 php to her father and 1500 was kept.

solution looks good.

solution is that she had 4000 php in the first place.